OpenStudy (anonymous):

factor the expression completely... 6x^3+48

OpenStudy (anonymous):

6x^3+48 First take 6 common

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

@mathgirl501 ???

OpenStudy (anonymous):

i don understand wht you are sking me @sauravshakya

OpenStudy (anonymous):

Can you take out a common number with 6 and 48?

OpenStudy (kira_yamato):

6x^3+48 = 6(x^3 + 8) = 6(x^3 + 2^3) = 6{(x + 2)[x^2 - (x)(2) + 2^2]} = 6(x + 2) (x^2 - 2x + 4)

OpenStudy (anonymous):

6x^3+48 = 6(x^3+8) = 6(x^3 +2^3) Now, USE a^3+b^3 = (a+b) (a^2-ab+b^2)

OpenStudy (anonymous):

Can u?